An irrigation canal contractor wants to determine whether he should purchase a used Caterpillar mini excavator, or a Toro powered rotary tiller for servicing irrigation ditches in an agricultural area of California. Caterpillar mini excavator: The initial cost of the excavator is $26,500 with a $9000 salvage value after 10 years. Fixed costs are expected to be $18,000 per year. The excavator will require maintenance and operating (M&O) cost of $107 per mile of ditch. Toro powered rotary tiller: Alternatively, the contractor can purchase a tiller. The tiller costs $1200 and has a useful life of 5 years with no salvage value. Its maintenance and operating (M&O) cost is expected to be $580 per mile of ditch. The contractor's MARR is 10% per year. A. Determine the number of miles of ditch per year the contractor would have to service for the two options to break even. B. If the number of miles of ditch per year is 60 miles which alternative is recommended? Caterpillar mini excavator Initial Cost $25,500 M&O Cost FC = $18,000/year VC = $107/mile Salvage Value $9,000 Service life 10 Toro powered rotary tiller Initial Cost $1,200 M&O Cost FC = $0/year VC = $580/mile Salvage Value Service life
Answer #1
B Present Value(PV) of cash flow
(Cash Flow)/((1+i)^N)
i=discount rate =MARR=10%=0.1
N=Year of Cash Flow
N Year 0 1 2 3 4 5 6 7 8 9 10
X Number of miles of ditch per year 60 60 60 60 60 60 60 60 60 60
Caterpillar Mini Excavator:
a Initial Cost $26,500
b Maintenance &Operating Cost per mile $107 $107 $107 $107 $107 $107 $107 $107 $107 $107
c=X*b Annual O&M Cost $6,420 $6,420 $6,420 $6,420 $6,420 $6,420 $6,420 $6,420 $6,420 $6,420
d Annual Fixed costs $18,000 $18,000 $18,000 $18,000 $18,000 $18,000 $18,000 $18,000 $18,000 $18,000
e Salvage Value ($9,000)
f=a+c+d+e Total Cash Outflow $26,500 $24,420 $24,420 $24,420 $24,420 $24,420 $24,420 $24,420 $24,420 $24,420 $15,420 SUM
PV1=f/(1.1^N) Present Value(PV) of Total cash flow $26,500 $22,200 $20,182 $18,347 $16,679 $15,163 $13,784 $12,531 $11,392 $10,356 $5,945 $173,080
NPW1 Net Present Worth of Costs=Sum of PV $173,080
Toro Powered Rotary Tiller:
g Initial Cost $1,200
h Maintenance &Operating Cost per mile $580 $580 $580 $580 $580
i=X*h Annual O&M Cost $34,800 $34,800 $34,800 $34,800 $34,800
j=g+i Cash Flow $1,200 $34,800 $34,800 $34,800 $34,800 $34,800
k Cash Flow for repeat investment $1,200 $34,800 $34,800 $34,800 $34,800 $34,800
l=j+k Total Cash Outflow $1,200 $34,800 $34,800 $34,800 $34,800 $36,000 $34,800 $34,800 $34,800 $34,800 $34,800 SUM
PV2=l/(1.1^N) Present Value(PV) of Total cash flow $1,200 $31,636 $28,760 $26,146 $23,769 $22,353 $19,644 $17,858 $16,234 $14,759 $13,417 $215,776
NPW2 Net Present Worth of Costs=Sum of PV $215,776
Net Present Worth of Costs is lower for Caterpillar Mini Excavator
Caterpillar Mini Excavator is recommended